I have this problem that I desperately need help with. I am supposed tro use factoring on this but I'm having trouble with it. It goes:
Tropical Pools sells an aboveground model for "p" dollars each. The monthly revenue from the sale of this product is given by
R = -0.08p^2 + 300p
Revenue is the product of "p" and the demand (quantity sold).
a.) Factor out the price on the right-hand side of the formula
b.) What is an expression of the monthly demand?
c.) What is the monthly demand for this pool when the price is $3000?
d.) What is the approximate maimum revenue?
Can anyone help me out with this?
Thanks

Answer by stanbon(48569)

The monthly revenue from the sale of this product is given by
R = -0.08p^2 + 300p
Revenue is the product of "p" and the demand (quantity sold).
a.) Factor out the price on the right-hand side of the formula
R=p(-0.08p+300)
b.) What is an expression of the monthly demand?
D=-0.08p+300
c.) What is the monthly demand for this pool when the price is $3000?
D(3000)=-0.08(3000)+300
D(3000)=240+300=540
d.) What is the approximate maximum revenue?
Maximum at -b/2a where a=-0.08, b=300
Max at -300/(-0.16)= 1875
R(1875)= -0.08(1875)^2+300(1875)=$281,250 (Maximum Revenue)
Cheers,
Stan H.